### v1

# Weak polynomial identities of small degree for the Weyl algebra

Abstract.
In this paper we investigate weak polynomial identities for the Weyl algebra $\mathsf{A}_{1}$ over an infinite field of arbitrary characteristic. Namely, we describe weak polynomial identities of the minimal degree, which is three, and of degrees 4 and 5. We also describe weak polynomial identities is two variables.
Keywords: polynomial identities, weak polynomial identities, matrix identities, Weyl algebra, positive characteristic.
2020 MSC: 16R10.
1. Introduction
Assume that $\mathbb{F}$ is an infinite field of arbitrary characteristic $p=\mathop{\rm char}\mathbb{F}\geq 0$. All vector spaces and algebras are over $\mathbb{F}$ and all algebras are associative, unless stated otherwise. We write $\mathbb{F}\langle x_{1},\ldots,x_{n}\rangle$ for the free unital $\mathbb{F}$-algebra with free generators $x_{1},\ldots,x_{n}$. In case the free generators are $x_{1},x_{2},\dotsc$ the corresponding free algebra is denoted by $\mathbb{F}\langle X\rangle$.
A polynomial identity for a unital $\mathbb{F}$-algebra $\mathcal{A}$ is an element $f(x_{1},\ldots,x_{m})$ of $\mathbb{F}\langle X\rangle$ such that $f(a_{1},\ldots,a_{m})=0$ in $\mathcal{A}$ for all $a_{1},\ldots,a_{m}\in\mathcal{A}$. The set ${{\rm Id}_{\mathbb{F}}({\mathcal{A}})}={{\rm Id}({\mathcal{A}})}$ of all polynomial identities for $\mathcal{A}$ is a T-ideal, i.e., ${{\rm Id}({\mathcal{A}})}$ is an ideal of $\mathbb{F}\langle X\rangle$ such that $\phi({{\rm Id}({\mathcal{A}})})\subset{{\rm Id}({\mathcal{A}})}$ for every endomorphism $\phi$ of $\mathbb{F}\langle X\rangle$. Given an $\mathbb{F}$-subspace $\mathcal{V}\subset\mathcal{A}$, we write ${{\rm Id}_{\mathbb{F}}({\mathcal{V}})}={{\rm Id}({\mathcal{V}})}$ for the ideal of all polynomial identities for $\mathcal{V}$. Note that ${{\rm Id}({\mathcal{V}})}$ is an L-ideal (or weak T-ideal), i.e., $\phi({{\rm Id}({\mathcal{V}})})\subset{{\rm Id}({\mathcal{V}})}$ for every linear endomorphism $\phi$ of $\mathbb{F}\langle X\rangle$, but ${{\rm Id}({\mathcal{V}})}$ is not a T-ideal in general. We say that a space $\mathcal{V}$ generates the algebra $\mathcal{A}$, if any element of $\mathcal{A}$ can be written as a non-commutative polynomial without free term in some elements of $\mathcal{V}$. If a space $\mathcal{V}$ generates the algebra $\mathcal{A}$, then the polynomial identities for $\mathcal{V}$ are called weak polynomial identities for the pair ($\mathcal{A},\mathcal{V}$) and we denote ${{\rm Id}({\mathcal{V}})}={{\rm Id}({\mathcal{A},\mathcal{V}})}$.
Weak polynomial identities were introduced in 1973 by Razmyslov [25, 26] (see also book [27]), who applied them to study polynomial identities of matrices. Razmyslov [25], Drensky [7] and Koshlukov [18] described weak polynomial identities for the pair $(M_{2},sl_{2})$ over a field of an arbitrary characteristic, where $sl_{2}$ is the space of all traceless matrices. Weak polynomial identities of small degrees for the pair $(M_{3},sl_{3})$ were studies by Drensky, Rashkova [6] and by Blachar, Matzri, Rowen, Vishne [2].
For $p=0$ weak polynomial identities for the pair $(M_{2},H_{2})$ were described by Drensky [4], where $H_{n}$ stands for the space of all symmetric $n\times n$ matrices. Minimal weak polynomial identities for the pair $(M_{n},H_{n})$ for an arbitrary $n>1$ were described by Ma and Racine [24] in case the characteristic of $\mathbb{F}$ satisfies certain restrictions.
Weak polynomial identities were also considered in [5, 12, 13, 17, 19], etc. More details on weak polynomial identities can be found in a recent survey by Drensky [8].
The Weyl algebra $\mathsf{A}_{1}$ is generated by $\mathsf{V}=\mathbb{F}\text{-}\mathrm{span}\{x,\;y\}$. In this paper we consider weak polynomial identities for the pair $(\mathsf{A}_{1},\mathsf{V})$. In Lemma 4.1 we show that the following elements of $\mathbb{F}\langle X\rangle$ are weak polynomial identities for $(\mathsf{A}_{1},\mathsf{V})$:
  1. $\bullet$
    $\Gamma_{m}(x_{1},\ldots,x_{m})=[[x_{1},x_{2}],x_{3}\cdots x_{m}]$ for $m\geq 3$,
  2. $\bullet$
    ${\rm St}_{3}(x_{1},x_{2},x_{3})=x_{1}[x_{2},x_{3}]-x_{2}[x_{1},x_{3}]+x_{3}[x_{1},x_{2}]$,
  3. $\bullet$
    $T_{4}(x_{1},\ldots,x_{4})=[x_{1},x_{2}][x_{3},x_{4}]-[x_{1},x_{3}][x_{2},x_{4}]+[x_{2},x_{3}][x_{1},x_{4}]$,
Denote by $\mathcal{I}$ the ideal of $\mathbb{F}\langle X\rangle$ generated by
$\Gamma_{3}(x_{i},x_{j},x_{k}),\;\;{\rm St}_{3}(x_{i},x_{j},x_{k}),\;\;T_{4}(x_{i},x_{j},x_{k},x_{l})$
for all $i,j,k,l>0$. In other words, $\mathcal{I}$ is the L-ideal generated by $\Gamma_{3}$, ${\rm St}_{3}$, and $T_{4}$. Given $f_{1},f_{2}\in\mathbb{F}\langle X\rangle$, we say that $f_{1}$ and $f_{2}$ are equivalent and write $f_{1}\equiv f_{2}$ in case $f_{1}-f_{2}\in\mathcal{I}$.
In Theorem 6.3 we describe weak polynomial identities for $(\mathsf{A}_{1},\mathsf{V})$ of the minimal degree, which is three. In Theorem 6.1 we show that every weak polynomial identity for $(\mathsf{A}_{1},\mathsf{V})$ in two variables lies in $\mathcal{I}$. Moreover, all weak polynomial identities for $(\mathsf{A}_{1},\mathsf{V})$ of degrees 4 and 5 belong to $\mathcal{I}$ by Propositions 7.1 and 7.2. Therefore, we formulate the following conjecture:
Conjecture 1.1.
The ideal of all weak polynomial identities for the pair $(\mathsf{A}_{1},\mathsf{V})$ is equal to $\mathcal{I}$.
The key definitions are given in Section 2 and some properties are considered in Section 3. The proofs are based on the notion of a completely reduced form of elements of $\mathbb{F}\langle X\rangle$, which is introduced in Section 5.
2. Definitions and known results
2.1. Polynomial identities for the Weyl algebra $\mathsf{A}_{1}$
The Weyl algebra $\mathsf{A}_{1}$ is the unital associative algebra over $\mathbb{F}$ generated by letters $x$, $y$ subject to the defining relation $yx=xy+1$ (equivalently, $[y,x]=1$, where $[y,x]=yx-xy$), i.e.,
$\mathsf{A}_{1}=\mathbb{F}\langle x,y\rangle/{{\rm id}\{{yx-xy-1}\}}.$
We say that algebras $\mathcal{A}$, $\mathcal{B}$ are called PI-equivalent and write $\mathcal{A}\sim_{\rm PI}\mathcal{B}$ if ${{\rm Id}({\mathcal{A}})}={{\rm Id}({\mathcal{B}})}$. We say that an L-ideal $I\in\mathbb{F}\langle X\rangle$ is generated by $f_{1},\ldots,f_{k}\in\mathbb{F}\langle X\rangle$ as an L-ideal, if $I$ is an $\mathbb{F}$-span of $\{f^{(1)}f_{i}(g_{1},\ldots,g_{m})f^{(2)}\}$ for all $f^{(1)},f^{(2)}\in\mathbb{F}\langle X\rangle$, all linear combinations $g_{1},\ldots,g_{m}$ of letters $\{x_{1},x_{2},\ldots\}$, and $1\leq i\leq k$. Obviously, in case $f_{i}$ is multilinear (see Section 2.2 below) we can assume that $g_{1},\ldots,g_{m}$ are letters.
Assume that $p=0$. It is well-known that the algebra $\mathsf{A}_{1}$ does not have nontrivial polynomial identities. Nevertheless, some subspaces of $\mathsf{A}_{1}$ satisfy certain polynomial identities. As an example, Dzhumadil’daev proved that the standard polynomial
${\rm St}_{N}(x_{1},\ldots,x_{N})=\sum_{\sigma\in\mathcal{S}_{N}}(-1)^{\sigma}x_{\sigma(1)}\cdots x_{\sigma(N)}$
is a polynomial identity for $\mathsf{A}_{1}^{(-,s)}=\mathbb{F}\text{-}\mathrm{span}\{ay^{s}\ |\ a\in\mathbb{F}[x]\}$ if and only if $N>2s$ (Theorem 1 of [10]). More results on polynomial identities for some subspaces of $n^{\rm th}$ Weyl algebra were obtained in [9, 11]. Considering $\mathsf{A}_{1}^{(-,1)}$ with respect to the Lie bracket we obtain is a simple Lie algebra $\mathsf{W}_{1}$, which is called Witt algebra. The well-known open conjecture claims that all polynomial identities for $\mathsf{W}_{1}$ follow from the standard Lie identity of degree 5. $\mathbb{Z}$-graded identities for $W_{1}$ were described by Freitas, Koshlukov and Krasilnikov [16]. Moreover, $\mathbb{Z}$-graded identities for the related Lie algebra of the derivations of the algebra of Laurent polynomials were described in [14, 15].
The situation is drastically different in case $p>0$. Namely, $\mathsf{A}_{1}$ is PI-equivalent to the algebra $M_{p}$ of all $p\times p$ matrices over $\mathbb{F}$. Moreover, the Weyl algebra $\mathsf{A}_{1}$ over an arbitrary associative (but possible non-commutative) $\mathbb{F}$-algebra $\mathsf{B}$ is PI-equivalent to the algebra $M_{p}(\mathsf{B})$ of all $p\times p$ matrices over $\mathsf{B}$ (see Theorem 4.9 of [21] for more general result). Polynomial identities for $\mathsf{A}_{1}^{(-,s)}$ and other subspaces of $\mathsf{A}_{1}$ were studied in [22, 23].
2.2. Notations
An algebra that satisfies a nontrivial polynomial identity is called a PI-algebra. A T-ideal $I$ of $\mathbb{F}\langle X\rangle$ generated by $f_{1},\ldots,f_{k}\in\mathbb{F}\langle X\rangle$ is the minimal T-ideal of $\mathbb{F}\langle X\rangle$ that contains $f_{1},\ldots,f_{k}$. We denote by $\langle X\rangle_{m}$ and $\langle X\rangle$ the monoids (with unity) freely generated by the letters $x_{1},\ldots,x_{m}$ and $x_{1},x_{2},\ldots$, respectively. Given $w\in\langle X\rangle_{m}$, we write $\deg_{x_{i}}(w)$ for the number of letters $x_{i}$ in $w$ and $\mathop{\rm mdeg}(w)\in\mathbb{N}_{0}^{m}$ for the multidegree $(\deg_{x_{1}}(w),\ldots,\deg_{x_{m}}(w))$ of $w$, where $\mathbb{N}_{0}=\{0,1,2,\ldots\}$ and =$\mathbb{N}=\{1,2,\ldots\}$. An element $f\in\mathbb{F}\langle X\rangle$ is called (multi)homogeneous if it is a linear combination of monomials of the same (multi)degree. Given $f=f(x_{1},\ldots,x_{m})$ of $\mathbb{F}\langle X\rangle$, we write $f=\sum_{{\underline{\delta}}\in\mathbb{N}_{0}^{m}}f_{{\underline{\delta}}}$ for multihomogeneous components $f_{{\underline{\delta}}}$ of $f$ with $\mathop{\rm mdeg}{f_{{\underline{\delta}}}}={\underline{\delta}}$. If $f\in\mathbb{F}\langle X\rangle$ is multihomogeneous of multidegree $1^{m}=(1,\ldots,1)$ ($m$ times), then $f$ is called multilinear. For ${\underline{\delta}}=(\delta_{1},\ldots,\delta_{m})$ we denote $|{\underline{\delta}}|=\delta_{1}+\cdots+\delta_{m}$. Given ${\underline{\delta}}\in\mathbb{N}_{0}^{m}$, we write $\mathbb{F}\langle X\rangle_{{\underline{\delta}}}$ for all elements of $\mathbb{F}\langle X\rangle$ of multidegree ${\underline{\delta}}$ and we write ${{\rm Id}({\mathcal{A},\mathcal{V}})}_{{\underline{\delta}}}$ for all elements of ${{\rm Id}({\mathcal{A},\mathcal{V}})}$ of multidegree ${\underline{\delta}}$.
3. Properties
3.1. Properties of $\mathsf{A}_{1}$
Given $a\in\mathbb{F}[x]$, we write $\partial(a)$ for the usual derivative of a polynomial $a$ with respect to the variable $x$. Using the linearity of derivative and induction on the degree of $a\in\mathbb{F}[x]$ it is easy to see that
$[y,a]=\partial(a)\text{ holds in }\mathsf{A}_{1}\text{ for all }a\in\mathbb{F}[x].$ (1)
The following properties are well-known (for example, see [1]):
Proposition 3.1.
  1. (a)
    $\{x^{i}y^{j}\ |\ i,j\geq 0\}$ and $\{y^{j}x^{i}\ |\ i,j\geq 0\}$ are $\mathbb{F}$-bases for $\mathsf{A}_{1}$.
  2. (b)
    If $p=0$, then the center ${\rm Z}(\mathsf{A}_{1})$ of $\mathsf{A}_{1}$ is $\mathbb{F}$; if $p>0$, then ${\rm Z}(\mathsf{A}_{1})=\mathbb{F}[x^{p},y^{p}]$.
  3. (c)
[…]
  4. (d)
    The algebra $\mathsf{A}_{1}$ is simple if and only if $p=0$.
3.2. Partial linearizations
Assume $f\in\mathbb{F}\langle X\rangle$ is multihomogeneous of multidegree ${\underline{\delta}}\in\mathbb{N}_{0}^{m}$. Given $1\leq i\leq m$ and ${\underline{\gamma}}\in\mathbb{N}_{0}^{k}$ for some $k>0$ with $|{\underline{\gamma}}|=\delta_{i}>0$, the partial linearization $\mathop{\rm lin}_{x_{i}}^{{\underline{\gamma}}}(f)$ of $f$ of multidegree ${\underline{\gamma}}$ with respect to $x_{i}$ is the multihomogeneous component of
$f(x_{1},\ldots,x_{i-1},x_{i}+\cdots+x_{i+k-1},x_{i+k},\ldots,x_{m+k-1})$
of multidegree $(\delta_{1},\ldots,\delta_{i-1},\gamma_{1},\ldots,\gamma_{k},\delta_{i+1},\ldots,\delta_{m}$). As an example,
$\rm{lin}_{x_{2}}^{(2,1)}(x_{1}^{2}x_{2}^{3}x_{3}^{2})=x_{1}^{2}(x_{2}^{2}x_{3}+x_{2}x_{3}x_{2}+x_{3}x_{2}^{2})x_{4}^{2}.$
The result of subsequent applications of partial linearizations to $f$ is also called a partial linearization of $f$. The complete linearization $\mathop{\rm lin}(f)$ of $f$ is the result of subsequent applications of $\mathop{\rm lin}_{x_{1}}^{1^{\delta_{1}}},\ldots,\mathop{\rm lin}_{x_{m}}^{1^{\delta_{m}}}$ to $f$.
Since $\mathbb{F}$ is infinite, it is well-known that the following lemma holds (see also Lemma 2.3 of [23]).
Lemma 3.2.
Assume $\mathcal{A}$ is a unital $\mathbb{F}$-algebra and $\mathcal{V}\subset\mathcal{A}$ is an $\mathbb{F}$-subspace.
  1. 1.
    If $f$ is a polynomial identity for $\mathcal{V}$, then all partial linearizations of $f$ are also polynomial identities for $\mathcal{V}$.
  2. 2.
    Assume that all partial linearizations of a multihomogeneous element $f$ of $\mathbb{F}\langle X\rangle$ are equal to zero over some basis of $\mathcal{V}$. Then $f$ is a polynomial identity for $\mathcal{V}$.
Note that part 1 of Lemma 3.2 does not hold in general for a finite field. As an example, see [20] for the case of $f(x_{1})=x_{1}^{n}$ and
$\mathcal{A}=\mathcal{V}=\frac{\mathbb{F}\langle X\rangle}{{{\rm id}\{{g^{n}\,|\,g\in\mathbb{F}\langle X\rangle\text{ without constant term}}\}}}.$
4. Identities
Lemma 4.1.
The following elements of $\mathbb{F}\langle X\rangle$ are weak polynomial identities for $(\mathsf{A}_{1},\mathsf{V})$:
  1. $\bullet$
    $\Gamma_{m}(x_{1},\ldots,x_{m})=[[x_{1},x_{2}],x_{3}\cdots x_{m}]$ for all $m\geq 3$,
  2. $\bullet$
    ${\rm St}_{3}(x_{1},x_{2},x_{3})=x_{1}[x_{2},x_{3}]-x_{2}[x_{1},x_{3}]+x_{3}[x_{1},x_{2}]$,
  3. $\bullet$
    $T_{4}(x_{1},\ldots,x_{4})=[x_{1},x_{2}][x_{3},x_{4}]-[x_{1},x_{3}][x_{2},x_{4}]+[x_{2},x_{3}][x_{1},x_{4}]$,
Proof.
1. Since $[x,x]=[y,y]=0$ and $[x,y]=-[y,x]=-1$, we have
$[u,v]\in Z(A_{1})\text{ for all }u,v\in\mathsf{V}.$ (2)
Thus $\Gamma_{m}\in{{\rm Id}({\mathsf{A}_{1},\mathsf{V}})}$.
Since any $g$ from the set $\{{\rm St}_{3},\;T_{4}\}$ is multilinear, to show that $g\in{{\rm Id}({\mathsf{A}_{1},\mathsf{V}})}$ it is enough to show that $g(u_{1},\ldots,u_{m})=0$ in $\mathsf{A}_{1}$ for all $u_{1},\ldots,u_{m}\in\{x,y\}$. Obviously, $g(x,\ldots,x)=g(y,\ldots,y)=0$.
2. Since ${\rm St}_{3}(x,x,y)={\rm St}_{3}(x,y,y)=0$, we obtain ${\rm St}_{3}\in{{\rm Id}({\mathsf{A}_{1},\mathsf{V}})}$.
3. If $(u_{1},\ldots,u_{4})$ is equal to $(x,x,x,y)$ or $(x,y,y,y)$, then $T_{4}(u_{\sigma(1)},\ldots,u_{\sigma(4)})=0$ for all $\sigma\in S_{4}$. Similarly to $T_{4}(x,x,y,y)=0-1+1=0$, we obtain that $T_{4}(u_{\sigma(1)},\ldots,u_{\sigma(4)})=0$ for all $\sigma\in S_{4}$ and $(u_{1},\ldots,u_{4})=(x,x,y,y)$. Thus $T_{4}\in{{\rm Id}({\mathsf{A}_{1},\mathsf{V}})}$. ∎
Lemma 4.2.
Any weak identity for the pair $(\mathsf{A}_{1},\mathsf{V})$ of degree $\leq 2$ is zero.
Proof.
By Lemma 3.2 it is enough to show that $f=0$ for every multihomogeneous $f\in{{\rm Id}({\mathsf{A}_{1},\mathsf{V}})}$ of degree $\leq 2$.
If $\mathop{\rm mdeg}(f)=(\delta)$ for $\delta\in\{1,2\}$, then $f=\alpha x_{1}^{\delta}$ for $\alpha\in\mathbb{F}$ and equality $f(x)=0$ implies $\alpha=0$.
Assume that $\mathop{\rm mdeg}(f)=(1,1)$ and $f=\alpha x_{1}x_{2}+\beta x_{2}x_{1}$ for $\alpha,\beta\in\mathbb{F}$. Then $f(x,x)=0$ implies that $\alpha+\beta=0$, i.e., $f=\alpha[x_{1},x_{2}]$. Hence $0=f(x,y)=-\alpha$ implies $\alpha=0$. ∎
Lemma 4.3.
The L-ideal generated by $\Gamma_{m}$, ${\rm St}_{3}$, $T_{3}$, where $m\geq 3$, coincides with $\mathcal{I}$.
Proof.
For $m>3$ consider
$\begin{array}[]{rcl}\Gamma_{m}(x_{1},\ldots,x_{m})&=&[x_{1},x_{2}]x_{3}x_{4}\cdots x_{m}-x_{3}x_{4}\cdots x_{m}[x_{1},x_{2}]\\ &\equiv&x_{3}[x_{1},x_{2}]x_{4}\cdots x_{m}-x_{3}x_{4}\cdots x_{m}[x_{1},x_{2}]\\ &\vdots&\\ &\equiv&x_{3}x_{4}\cdots x_{m}[x_{1},x_{2}]-x_{3}x_{4}\cdots x_{m}[x_{1},x_{2}]=0.\\ \end{array}$
The claim is proven. ∎
5. Completely reduced bracket-monomials
Definition 5.1.
A product
$x_{t_{1}}\cdots x_{t_{l}}[x_{r_{1}},x_{s_{1}}]\cdots[x_{r_{k}},x_{s_{k}}]$
from $\mathbb{F}\langle X\rangle$, where ${\underline{t}}\in\mathbb{N}^{l}$, ${\underline{r}},{\underline{s}}\in\mathbb{N}^{k}$ for some $l\geq 0$, $k>0$ with $r_{1}<s_{1},\ldots,r_{k}<s_{k}$, is called a bracket-monomial.
Lemma 5.2.
If two bracket-monomials are equal in $\mathbb{F}\langle X\rangle$, then they are the same. In other words, if
$f=x_{t_{1}}\cdots x_{t_{l}}[x_{r_{1}},x_{s_{1}}]\cdots[x_{r_{k}},x_{s_{k}}]\;\text{ and }\;f^{\prime}=x_{t^{\prime}_{1}}\ldots x_{t^{\prime}_{l^{\prime}}}[x_{r^{\prime}_{1}},x_{s^{\prime}_{1}}]\cdots[x_{r^{\prime}_{k^{\prime}}},x_{s^{\prime}_{k^{\prime}}}]$
are bracket-monomials and $f=f^{\prime}$ in $\mathbb{F}\langle X\rangle$, then ${\underline{t}}={\underline{t}}^{\prime}$, ${\underline{r}}={\underline{r}}^{\prime}$, ${\underline{s}}={\underline{s}}^{\prime}$.
Proof.
Note that $f$ can be written as a linear combinations of $2^{k}$ pairwise different monomials from $\langle X\rangle$ with coefficients ±$\pm 1$.
We can assume that $l\geq l^{\prime}$. Thus, $t_{1}=t^{\prime}_{1},\ldots,t_{l^{\prime}}=t^{\prime}_{l^{\prime}}$. Therefore, without loss of generality, we may assume that $l^{\prime}=0$. In case $l>0$ we obtain that $f$ is a linear combination of pairwise different monomials which start with $x_{t_{1}}$, but $f^{\prime}$ is a linear combination of pairwise different monomials which start with $x_{r^{\prime}_{1}}$ and $x_{s^{\prime}_{1}}$, where $r^{\prime}_{1}\neq s^{\prime}_{1}$; a contradiction. Therefore, $l^{\prime}=0$.
We have that $f$ is a linear combination of pairwise different monomials which start with $x_{r_{1}}$ and $x_{s_{1}}$, but $f^{\prime}$ is a linear combination of pairwise different monomials which start with $x_{r^{\prime}_{1}}$ and $x_{s^{\prime}_{1}}$. Hence, $\{r_{1},s_{1}\}=\{r^{\prime}_{1},s^{\prime}_{1}\}$ and inequalities $r_{1}<s_{1}$, $r^{\prime}_{1}<s^{\prime}_{1}$ imply that $r_{1}=r^{\prime}_{1}$ and $s_{1}=s^{\prime}_{1}$. Therefore, without loss of generality we can assume that $f=[x_{r_{2}},x_{s_{2}}]\cdots[x_{r_{k}},x_{s_{k}}]$ and $f^{\prime}=[x_{r^{\prime}_{2}},x_{s^{\prime}_{2}}]\cdots[x_{r^{\prime}_{k^{\prime}}},x_{s^{\prime}_{k^{\prime}}}]$. Repeating the above reasoning several times we conclude the proof. ∎
Definition 5.3.
(a) A bracket-monomial
$f=x_{t_{1}}\cdots x_{t_{l}}[x_{r_{1}},x_{s_{1}}]\cdots[x_{r_{k}},x_{s_{k}}]\in\mathbb{F}\langle X\rangle,$ (3)
where ${\underline{t}}\in\mathbb{N}^{l}$, ${\underline{r}},{\underline{s}}\in\mathbb{N}^{k}$ for some $l\geq 0$, $k>0$ with $r_{1}<s_{1},\ldots,r_{k}<s_{k}$ is semi-reduced if
  1. $\bullet$
    $t_{1}\leq\cdots\leq t_{l}$;
  2. $\bullet$
    $s_{1}\leq\cdots\leq s_{k}$.
(b) A semi-reduced bracket-monomial $f\in\mathbb{F}\langle X\rangle$ defined by (3) is reduced if
  1. $\bullet$
    either $l=0$ or $l\geq 1$, $t_{l}\leq s_{1}$.
(c) A reduced bracket-monomial $f\in\mathbb{F}\langle X\rangle$ defined by (3) is completely reduced if
  1. $\bullet$
    do not exist $1\leq i\neq j\leq k$ with $r_{j}<r_{i}<s_{i}<s_{j}$.
Example 5.4.
Consider is the list of all completely reduced bracket-monomials of multidegree $1^{m}$:
  1. $m=2$:
    $[x_{1},x_{2}]$;
  2. $m=3$:
    $x_{1}[x_{2},x_{3}]$, $x_{2}[x_{1},x_{3}]$;
  3. $m=4$:
    $x_{1}x_{2}[x_{3},x_{4}]$, $x_{1}x_{3}[x_{2},x_{4}]$, $x_{2}x_{3}[x_{1},x_{4}]$, $[x_{1},x_{2}][x_{3},x_{4}]$, $[x_{1},x_{3}][x_{2},x_{4}]$;
  4. $m=5$:
    $x_{1}x_{2}x_{3}[x_{4},x_{5}]$, $x_{1}x_{2}x_{4}[x_{3},x_{5}]$, $x_{1}x_{3}x_{4}[x_{2},x_{5}]$, $x_{2}x_{3}x_{4}[x_{1},x_{5}]$,
    $x_{1}[x_{2},x_{3}][x_{4},x_{5}]$, $x_{1}[x_{2},x_{4}][x_{3},x_{5}]$, $x_{2}[x_{1},x_{3}][x_{4},x_{5}]$, $x_{2}[x_{1},x_{4}][x_{3},x_{5}]$,
    $x_{3}[x_{1},x_{4}][x_{2},x_{5}]$.
For $1\leq i<j$ we consider $\mathbb{N}_{0}^{i}$ as a subset of $\mathbb{N}_{0}^{j}$ by
$(r_{1},\ldots,r_{i})\to(r_{1},\ldots,r_{i},\underbrace{0,\ldots,0}_{j-i}).$
Assume ${\underline{r}}\in\mathbb{N}_{0}^{i}$ and ${\underline{s}}\in\mathbb{N}_{0}^{j}$ for some $i,j\geq 1$. Then we write ${\underline{r}}<{\underline{s}}$ for the lexicographical order on $\mathbb{N}_{0}^{k}$, where $k=\max\{i,j\}$ and we consider ${\underline{r}}$, ${\underline{s}}$ as elements of $\mathbb{N}_{0}^{k}$.
Definition 5.5.
Consider a bracket-monomial
$f=x_{t_{1}}\cdots x_{t_{l}}[x_{r_{1}},x_{s_{1}}]\cdots[x_{r_{k}},x_{s_{k}}]\in\mathbb{F}\langle X\rangle,$ (4)
where ${\underline{t}}\in\mathbb{N}^{l}$, ${\underline{r}},{\underline{s}}\in\mathbb{N}^{k}$ for some $l\geq 0$, $k>0$ and $r_{1}<s_{1},\ldots,r_{k}<s_{k}$. Then
  1. (a)
    the monomial weight of $f$ in case $l>0$ is $\mathop{\rm mw}(f)=(t_{\sigma(1)},\ldots,t_{\sigma(l)})$ for some permutation $\sigma\in\mathcal{S}_{l}$ such that
    $t_{\sigma(1)}\geq\cdots\geq t_{\sigma(l)}$
    and $\mathop{\rm mw}(f)=(0)$ in case $l=0$.
  2. (b)
    the bracket weight of $f$ is $\mathop{\rm bw}(f)=(s_{\sigma(1)}-r_{\sigma(1)},\ldots,s_{\sigma(k)}-r_{\sigma(k)})$ for some permutation $\sigma\in\mathcal{S}_{k}$ such that
    $s_{\sigma(1)}-r_{\sigma(1)}\geq\cdots\geq s_{\sigma(k)}-r_{\sigma(k)}.$
Example 5.6.
(a) We have ${\rm St}_{3}(x_{1},x_{2},x_{3})=f_{1}-f_{2}+f_{3}$ for the semi-reduced bracket-monomials
$f_{1}=x_{1}[x_{2},x_{3}],\;\;f_{2}=x_{2}[x_{1},x_{3}],\;\;f_{3}=x_{3}[x_{1},x_{2}].$
Then $\mathop{\rm mw}(f_{1})=(1)$, $\mathop{\rm mw}(f_{2})=(2)$, $\mathop{\rm mw}(f_{3})=(3)$ and $\mathop{\rm bw}(f_{1})=(1)$, $\mathop{\rm bw}(f_{2})=(2)$, $\mathop{\rm bw}(f_{3})=(1)$. Note that $f_{1}$, $f_{2}$ are reduced, but $f_{3}$ is not reduced.
(b) We have $T_{4}(x_{1},\ldots,x_{4})=h_{1}-h_{2}+h_{3}$ for the reduced bracket-monomials
$h_{1}=[x_{1},x_{2}][x_{3},x_{4}],\;\;h_{2}=[x_{1},x_{3}][x_{2},x_{4}],\;\;h_{3}=[x_{2},x_{3}][x_{1},x_{4}].$
Then $\mathop{\rm mw}(h_{i})=(0)$ for $i=1,2,3$ and $\mathop{\rm bw}(h_{1})=(1,1)$, $\mathop{\rm bw}(h_{2})=(2,2)$, $\mathop{\rm bw}(h_{3})=(3,1)$. Note that $h_{1}$, $h_{2}$ are completely reduced, but $h_{3}$ is not completely reduced.
Lemma 5.7.
Assume that $f\in\mathbb{F}\langle X\rangle$ is multihomogeneous of multidegree ${\underline{\delta}}\in\mathbb{N}_{0}^{m}$. Then there are semi-reduced bracket-monomials $f_{i}\in\mathbb{F}\langle X\rangle$ and $\alpha_{i},\beta\in\mathbb{F}$ such that
$f\equiv\beta x_{1}^{\delta_{1}}\cdots x_{m}^{\delta_{m}}+\sum_{i}\alpha_{i}f_{i},$
where $\mathop{\rm mdeg}(f_{i})={\underline{\delta}}$ for all $i$. Moreover,
  1. (a)
    if $f\in{{\rm Id}({\mathsf{A}_{1},\mathsf{V}})}$, then $\beta=0$;
  2. (b)
    if $f$ is a bracket-monomial, then $\beta=0$ and $\mathop{\rm mw}(f_{i})\leq\mathop{\rm mw}(f)$ for all $i$.
Proof.
Assume that $f_{1},f_{2}\in\mathbb{F}\langle x_{1},\ldots,x_{m}\rangle$ are multihomogeneous and $1\leq i<j\leq m$. Since
$f_{1}x_{j}x_{i}f_{2}=f_{1}x_{i}x_{j}f_{2}-f_{1}[x_{i},x_{j}]f_{2}=f_{1}x_{i}x_{j}f_{2}-f_{1}f_{2}[x_{i},x_{j}]-f_{1}[[x_{i},x_{j}],f_{2}],$
by Lemma 4.3 we obtain that
$f_{1}x_{j}x_{i}f_{2}\equiv f_{1}x_{i}x_{j}f_{2}-f_{1}f_{2}[x_{i},x_{j}].$ (5)
Lemma 4.3 also implies that
$f_{1}[x_{i},x_{j}]f_{0}f_{2}\equiv f_{1}f_{0}[x_{i},x_{j}]f_{2}$ (6)
for every $f_{0}\in\mathbb{F}\langle X\rangle$. Since equivalences (5), (6) preserve the multidegree, applying formulas (5), (6) to $f$ we obtain that $f\equiv\beta x_{1}^{\delta_{1}}\cdots x_{m}^{\delta_{m}}+\sum_{i}\alpha_{i}f_{i}$ for some semi-reduced bracket-monomials $f_{i}$ and $\beta,\alpha_{i}\in\mathbb{F}$, where $\mathop{\rm mdeg}(f_{i})={\underline{\delta}}$.
Assume $f\in{{\rm Id}({\mathsf{A}_{1},\mathsf{V}})}$. Since $[x,x]=0$, we have $f_{i}(x,\ldots,x)=0$ in $\mathsf{A}_{1}$ for all $i$. Therefore, $0=f(x,\ldots,x)=\beta\,x^{|{\underline{\delta}}|}$. Thus $\beta=0$.
If $f$ is a bracket-monomial, then it is easy to see that $\beta=0$ and $\mathop{\rm mw}(f_{i})\leq\mathop{\rm mw}(f)$ for all $i$. ∎
Lemma 5.8.
Consider a semi-reduced bracket-monomial
$f=x_{t_{1}}\cdots x_{t_{l}}[x_{r_{1}},x_{s_{1}}]\cdots[x_{r_{k}},x_{s_{k}}]\in\mathbb{F}\langle X\rangle.$
Then $f\equiv\sum_{i}\alpha_{i}f_{i}$ for some $\alpha_{i}\in\mathbb{F}$, $f_{i}\in\mathbb{F}\langle X\rangle$ such that $f_{i}$ is a reduced bracket-monomial, $\mathop{\rm mdeg}(f_{i})=\mathop{\rm mdeg}(f)$, and $\mathop{\rm mw}(f_{i})\leq\mathop{\rm mw}(f)$ for all $i$.
Proof.
Since $f$ is semi-reduced, we have $r_{1}<s_{1},\ldots,r_{k}<s_{k}$, $t_{1}\leq\cdots\leq t_{l}$ and $s_{1}\leq\cdots\leq s_{k}$, where $l\geq 0$, $k>0$.
We prove the lemma by induction on $\mathop{\rm mw}(f)$. If $\mathop{\rm mw}(f)=(0)$, then $l=0$ and $f$ is reduced.
Assume that $(0)<\mathop{\rm mw}(f)$ and for every semi-reduced bracket-monomial $f^{\prime}\in\mathbb{F}\langle X\rangle$ with $\mathop{\rm mw}(f^{\prime})<\mathop{\rm mw}(f)$ the statement of this lemma holds.
Assume that $f$ is not reduced, i.e., $l\geq 1$ and $t_{l}>s_{1}$. Using ${\rm St}_{3}$ from Lemma 4.1, we obtain
$x_{t_{l}}[x_{r_{1}},x_{s_{1}}]\equiv x_{s_{1}}[x_{r_{1}},x_{t_{l}}]-x_{r_{1}}[x_{s_{1}},x_{t_{l}}].$ (7)
Thus $f\equiv f_{1}-f_{2}$ for
$f_{1}=x_{t_{1}}\cdots x_{t_{l-1}}x_{s_{1}}[x_{r_{1}},x_{t_{l}}][x_{r_{2}},x_{s_{2}}]\cdots[x_{r_{k}},x_{s_{k}}],$
$f_{2}=x_{t_{1}}\cdots x_{t_{l-1}}x_{r_{1}}[x_{s_{1}},x_{t_{l}}][x_{r_{2}},x_{s_{2}}]\cdots[x_{r_{k}},x_{s_{k}}].$
Note that $\mathop{\rm mdeg}(f_{1})=\mathop{\rm mdeg}(f_{2})=\mathop{\rm mdeg}(f)$. Using part (b) of Lemma 5.7, we obtain semi-reduced bracket-monomials $g_{1i},g_{2j}\in\mathbb{F}\langle X\rangle$ and $\alpha_{1i},\alpha_{2j}\in\mathbb{F}$ such that $f_{1}\equiv\sum_{i}\alpha_{1i}g_{1i}$ and $f_{2}\equiv\sum_{j}\alpha_{2j}g_{2j}$, where $\mathop{\rm mdeg}(g_{1i})=\mathop{\rm mdeg}(g_{2j})=\mathop{\rm mdeg}(f)$, $\mathop{\rm mw}(g_{1i})\leq\mathop{\rm mw}(f_{1})<\mathop{\rm mw}(f)$ and $\mathop{\rm mw}(g_{2j})\leq\mathop{\rm mw}(f_{2})<\mathop{\rm mw}(f)$. Applying the induction hypothesis to $g_{1i}$, $g_{2j}$ we conclude the proof, since $f\equiv\sum_{i}\alpha_{1i}g_{1i}-\sum_{i}\alpha_{2i}g_{2i}$. ∎
Lemma 5.9.
Consider a reduced bracket-monomial
$f=x_{t_{1}}\cdots x_{t_{l}}[x_{r_{1}},x_{s_{1}}]\cdots[x_{r_{k}},x_{s_{k}}]\in\mathbb{F}\langle X\rangle.$
Then $f\equiv\sum_{i}\alpha_{i}f_{i}$ for some $\alpha_{i}\in\mathbb{F}$, $f_{i}\in\mathbb{F}\langle X\rangle$ such that $f_{i}$ is a completely reduced bracket-monomial, $\mathop{\rm mdeg}(f_{i})=\mathop{\rm mdeg}(f)$, and $\mathop{\rm mw}(f_{i})\leq\mathop{\rm mw}(f)$ for all $i$.
Proof.
Since $f$ is reduced, we have $r_{1}<s_{1},\ldots,r_{k}<s_{k}$, $t_{1}\leq\cdots\leq t_{l}\leq s_{1}\leq\cdots\leq s_{k}$, where $l\geq 0$, $k>0$.
We prove the lemma by induction on $\mathop{\rm mw}(f)$.
1. Assume $\mathop{\rm mw}(f)=(0)$, i.e., $f=[x_{r_{1}},x_{s_{1}}]\cdots[x_{r_{k}},x_{s_{k}}]$. To show that the statement of this lemma holds for $f$ we use induction on $\mathop{\rm bw}(f)$.
Obviously, if $\mathop{\rm bw}(f)=(1)$, i.e., $f=[x_{r_{1}},x_{s_{1}}]$ with $s_{1}-r_{1}=1$, then $f$ is completely reduced.
Assume $(1)<\mathop{\rm bw}(f)$ and for every reduced bracket-monomial $f^{\prime}\in\mathbb{F}\langle X\rangle$ with $\mathop{\rm mw}(f^{\prime})=(0)$ and $\mathop{\rm bw}(f^{\prime})<\mathop{\rm bw}(f)$ the statement of the lemma holds.
Assume that $f$ is not completely reduced, i.e., there are $1\leq i\neq j\leq k$ such that $r_{j}<r_{i}<s_{i}<s_{j}$. For short, denote $a_{1}=r_{j}$, $a_{2}=r_{i}$, $a_{3}=s_{i}$, $a_{4}=s_{j}$. Note that $a_{1}<a_{2}<a_{3}<a_{4}$. Using equivalence (6) we obtain that
$[x_{r_{1}},x_{s_{1}}]\cdots[x_{r_{k}},x_{s_{k}}]\equiv[x_{a_{2}},x_{a_{3}}][x_{a_{1}},x_{a_{4}}]\,b$
for some product $b=[x_{r^{\prime}_{1}},x_{s^{\prime}_{1}}]\cdots[x_{r^{\prime}_{k-2}},x_{s^{\prime}_{k-2}}]$ of brackets. Applying equivalence $T_{4}(x_{a_{1}},x_{a_{2}},x_{a_{3}},x_{a_{4}})\equiv 0$, we obtain
$[x_{a_{2}},x_{a_{3}}][x_{a_{1}},x_{a_{4}}]\equiv-[x_{a_{1}},x_{a_{2}}][x_{a_{3}},x_{a_{4}}]+[x_{a_{1}},x_{a_{3}}][x_{a_{2}},x_{a_{4}}].$
Thus $f\equiv-f_{1}+f_{2}$ for
$f_{1}=[x_{a_{1}},x_{a_{2}}][x_{a_{3}},x_{a_{4}}]b,$
$f_{2}=[x_{a_{1}},x_{a_{3}}][x_{a_{2}},x_{a_{4}}]b.$
Note that $\mathop{\rm mdeg}(f_{1})=\mathop{\rm mdeg}(f_{2})=\mathop{\rm mdeg}(f)$. Since
$\mathop{\rm bw}([x_{a_{2}},x_{a_{3}}][x_{a_{1}},x_{a_{4}}])=(a_{4}-a_{1},a_{3}-a_{2})$
is greater than $\mathop{\rm bw}([x_{a_{1}},x_{a_{2}}][x_{a_{3}},x_{a_{4}}])$ and $\mathop{\rm bw}([x_{a_{1}},x_{a_{3}}][x_{a_{2}},x_{a_{4}}])$, we can see that $\mathop{\rm bw}(f_{1})<\mathop{\rm bw}(f)$ and $\mathop{\rm bw}(f_{2})<\mathop{\rm bw}(f)$.
We use equivalence (6) to obtain reduced bracket-monomials $g_{1}$ and $g_{2}$ such that $g_{1}\equiv f_{1}$ and $g_{2}\equiv f_{2}$, where $\mathop{\rm mdeg}(g_{1})=\mathop{\rm mdeg}(g_{2})=\mathop{\rm mdeg}(f)$, $\mathop{\rm mw}(g_{1})=\mathop{\rm mw}(g_{2})=(0)$,
$\mathop{\rm bw}(g_{1})=\mathop{\rm bw}(f_{1})<\mathop{\rm bw}(f)\;\;\text{ and}\;\;\mathop{\rm bw}(g_{2})=\mathop{\rm bw}(f_{2})<\mathop{\rm bw}(f).$
Applying the induction hypothesis to $g_{1}$ and $g_{2}$, we can see that the statement of this lemma holds for $f$.
2. Assume that $(0)<\mathop{\rm mw}(f)$, i.e., $l\geq 1$, and for every reduced bracket-monomial $f^{\prime}\in\mathbb{F}\langle X\rangle$ with $\mathop{\rm mw}(f^{\prime})<\mathop{\rm mw}(f)$ the claim of this lemma holds. To show that the statement of this lemma holds for $f$ we use induction on $\mathop{\rm bw}(f)$.
Obviously, if $\mathop{\rm bw}(f)=(1)$, i.e., $f=x_{t_{1}}\cdots x_{t_{l}}[x_{r_{1}},x_{s_{1}}]$ with $s_{1}-r_{1}=1$, then $f$ is completely reduced.
Assume $(1)<\mathop{\rm bw}(f)$ and for every reduced bracket-monomial $f^{\prime}\in\mathbb{F}\langle X\rangle$ with $\mathop{\rm mw}(f^{\prime})=\mathop{\rm mw}(f)$ and $\mathop{\rm bw}(f^{\prime})<\mathop{\rm bw}(f)$ the statement of the lemma holds.
Assume that $f$ is not completely reduced, i.e., there are $1\leq i\neq j\leq k$ such that $r_{j}<r_{i}<s_{i}<s_{j}$. For short, denote $a_{1}=r_{j}$, $a_{2}=r_{i}$, $a_{3}=s_{i}$, $a_{4}=s_{j}$. Note that $a_{1}<a_{2}<a_{3}<a_{4}$ and $t_{l}\leq a_{3}$. Using equivalence (6) we obtain that
$[x_{r_{1}},x_{s_{1}}]\cdots[x_{r_{k}},x_{s_{k}}]\equiv[x_{a_{2}},x_{a_{3}}][x_{a_{1}},x_{a_{4}}]\,b$
for some product $b=[x_{r^{\prime}_{1}},x_{s^{\prime}_{1}}]\cdots[x_{r^{\prime}_{k-2}},x_{s^{\prime}_{k-2}}]$ of brackets, where $t_{l}\leq s^{\prime}_{1}\leq\cdots\leq s^{\prime}_{k-2}$. Since $T_{4}(x_{a_{1}},x_{a_{2}},x_{a_{3}},x_{a_{4}})\equiv 0$, we obtain
$[x_{a_{2}},x_{a_{3}}][x_{a_{1}},x_{a_{4}}]\equiv-[x_{a_{1}},x_{a_{2}}][x_{a_{3}},x_{a_{4}}]+[x_{a_{1}},x_{a_{3}}][x_{a_{2}},x_{a_{4}}].$
Thus $f\equiv-f_{1}+f_{2}$ for
$f_{1}=x_{t_{1}}\cdots x_{t_{l}}[x_{a_{1}},x_{a_{2}}][x_{a_{3}},x_{a_{4}}]b,$
$f_{2}=x_{t_{1}}\cdots x_{t_{l}}[x_{a_{1}},x_{a_{3}}][x_{a_{2}},x_{a_{4}}]b.$
Note that $\mathop{\rm mdeg}(f_{1})=\mathop{\rm mdeg}(f_{2})=\mathop{\rm mdeg}(f)$. Since
$\mathop{\rm bw}([x_{a_{2}},x_{a_{3}}][x_{a_{1}},x_{a_{4}}])=(a_{4}-a_{1},a_{3}-a_{2})$
is greater than $\mathop{\rm bw}([x_{a_{1}},x_{a_{2}}][x_{a_{3}},x_{a_{4}}])$ and $\mathop{\rm bw}([x_{a_{1}},x_{a_{3}}][x_{a_{2}},x_{a_{4}}])$, we can see that $\mathop{\rm bw}(f_{1})<\mathop{\rm bw}(f)$ and $\mathop{\rm bw}(f_{2})<\mathop{\rm bw}(f)$.
2.1. Assume that $l_{t}\leq a_{2}$. Then $l_{t}\leq a_{2}<a_{3}<a_{4}$ and $t_{l}\leq s^{\prime}_{1}\leq\cdots\leq s^{\prime}_{k-2}$. We use equivalence (6) to obtain reduced $g_{1}$ and $g_{2}$ such that $g_{1}\equiv f_{1}$ and $g_{2}\equiv f_{2}$, where $\mathop{\rm mdeg}(g_{1})=\mathop{\rm mdeg}(g_{2})=\mathop{\rm mdeg}(f)$, $\mathop{\rm mw}(g_{1})=\mathop{\rm mw}(g_{2})=\mathop{\rm mw}(f)$,
$\mathop{\rm bw}(g_{1})=\mathop{\rm bw}(f_{1})<\mathop{\rm bw}(f)\;\;\text{ and}\;\;\mathop{\rm bw}(g_{2})=\mathop{\rm bw}(f_{2})<\mathop{\rm bw}(f).$
Applying induction on brace weight to $g_{1}$ and $g_{2}$, we obtain that the statement of the lemma holds for $f$.
2.2. Assume $a_{2}<t_{l}$. Using equivalence (7), we obtain $f_{1}\equiv h_{1}-h_{2}$ for
$h_{1}=x_{t_{1}}\cdots x_{t_{l-1}}x_{a_{2}}[x_{a_{1}},t_{l}][x_{a_{3}},x_{a_{4}}]b,$
$h_{2}=x_{t_{1}}\cdots x_{t_{l-1}}x_{a_{1}}[x_{a_{2}},t_{l}][x_{a_{3}},x_{a_{4}}]b.$
Since $t_{l}>a_{1},a_{2}$, we have $\mathop{\rm mw}(h_{1})<\mathop{\rm mw}(f)$ and $\mathop{\rm mw}(h_{2})<\mathop{\rm mw}(f)$. Using part (b) of Lemma 5.7 and Lemma 5.8, we obtain reduced bracket-monomials $g_{1i^{\prime}},g_{2j^{\prime}}\in\mathbb{F}\langle X\rangle$ and $\alpha_{1i^{\prime}},\alpha_{2j^{\prime}}\in\mathbb{F}$ such that $h_{1}\equiv\sum_{i^{\prime}}\alpha_{1i^{\prime}}g_{1i^{\prime}}$ and $h_{2}\equiv\sum_{j^{\prime}}\alpha_{2j^{\prime}}g_{2j^{\prime}}$, where $\mathop{\rm mdeg}(g_{1i^{\prime}})=\mathop{\rm mdeg}(g_{2j^{\prime}})=\mathop{\rm mdeg}(f)$,
$\mathop{\rm mw}(g_{1i^{\prime}})\leq\mathop{\rm mw}(h_{1})<\mathop{\rm mw}(f)\;\;\text{ and }\;\;\mathop{\rm mw}(g_{2j^{\prime}})\leq\mathop{\rm mw}(h_{2})<\mathop{\rm mw}(f).$
We apply induction on monomial weight to $g_{1i^{\prime}}$ and $g_{2j^{\prime}}$ to show that the statement of the lemma holds for $f_{1}$. We establish that the statement of the lemma holds for $f_{2}$ by repeating the proof from part 2.1. Therefore, the statement of the lemma holds for $f$. ∎
Theorem 5.10.
Assume that $f\in\mathbb{F}\langle X\rangle$ is multihomogeneous of multidegree ${\underline{\delta}}\in\mathbb{N}_{0}^{m}$. Then there are completely reduced bracket-monomials $f_{i}\in\mathbb{F}\langle X\rangle$ and $\alpha_{i},\beta\in\mathbb{F}$ such that
$f\equiv\beta x_{1}^{\delta_{1}}\cdots x_{m}^{\delta_{m}}+\sum_{i}\alpha_{i}f_{i},$
where $\mathop{\rm mdeg}(f_{i})={\underline{\delta}}$ for all $i$. Moreover,
  1. (a)
    if $f\in{{\rm Id}({\mathsf{A}_{1},\mathsf{V}})}$, then $\beta=0$;
  2. (b)
    if $f$ is a bracket-monomial, then $\beta=0$ and $\mathop{\rm mw}(f_{i})\leq\mathop{\rm mw}(f)$ for all $i$.
Proof.
Consequently applying Lemmas 5.7, 5.8, 5.9 we obtain the required. ∎
6. Minimal weak polynomial identities
Theorem 6.1.
Every weak polynomial identity for the pair $(\mathsf{A}_{1},\mathsf{V})$ in two variables lies in the L-ideal $\mathcal{I}$ generated by ${\rm St}_{3}$, $\Gamma_{3}$, $T_{4}$.
Proof.
Assume that $f\in\mathbb{F}\langle x_{1},x_{2}\rangle$ is a weak polynomial identity in two variables for the pair $(\mathsf{A}_{1},\mathsf{V})$. By Lemma 3.2, we can assume that $f$ is multihomogeneous of multidegree $(r,s)$ for some $r,s\geq 0$. Then Theorem 5.10 implies that $f$ is equivalent to a linear combination of completely reduced bracket-monomials of multidegree $(r,s)$.
Assume that $r\geq s$. If $s=0$, then $f\equiv 0$ by the definition of completely reduced bracket-monomials. Assume $s>0$. Then
$f(x_{1},x_{2})\equiv\sum_{i=1}^{s}\alpha_{i}x_{1}^{r-i}x_{2}^{s-i}[x_{1},x_{2}]^{i}$
for some $\alpha_{i}\in\mathbb{F}$. Since $0=f(x,y)=\sum_{i=1}^{s}(-1)^{i}\alpha_{i}x^{r-i}y^{s-i}$ in $\mathsf{A}_{1}$, we obtain that $\alpha_{1}=\cdots=\alpha_{s}=0$ by part (a) of Proposition 3.1, i.e., $f\equiv 0$.
The case of $r<s$ can be considered similarly. The proof is completed. ∎
Lemma 6.2.
Every weak polynomial identity for the pair $(\mathsf{A}_{1},\mathsf{V})$ of degree $3$ lies in the L-ideal generated by $\Gamma_{3}$ and ${\rm St}_{3}$.
Proof.
Assume that $f\in\mathbb{F}\langle X\rangle$ is a weak polynomial identity of degree 3 for the pair $(\mathsf{A}_{1},\mathsf{V})$. By Lemma 3.2, we can assume that $f$ is multihomogeneous of multidegree $\Delta$ with $|\Delta|=3$. By Theorem 5.10, $f$ is equivalent to a linear combination of completely reduced bracket-monomials of multidegree $\Delta$.
Assume $\Delta=(1,1,1)$. Then
$f(x_{1},x_{2},x_{3})\equiv\alpha x_{1}[x_{2},x_{3}]+\beta x_{2}[x_{1},x_{3}],$
where $\alpha,\beta\in\mathbb{F}$. Since $0=f(x,y,y)=-\beta y$ and $0=f(x,y,x)=\alpha x$ in $\mathsf{A}_{1}$, we obtain $\alpha=\beta=0$. The definition of the ideal $\mathcal{I}$ implies the required.
If $\Delta=(2,1)$ or $\Delta=(3)$, then Theorem 6.1 concludes the proof. ∎
Theorem 6.3.
1. The minimal degree of a non-trivial weak polynomial identity for the pair $(\mathsf{A}_{1},\mathsf{V})$ is three.
2. The vector space ${{\rm Id}({\mathsf{A}_{1},\mathsf{V}})}_{\Delta}$ for $|\Delta|=3$ has the following basis:
  1. $\bullet$
    $\Gamma_{3}(x_{1},x_{2},x_{3})$, $\Gamma_{3}(x_{1},x_{3},x_{2})$, and ${\rm St}_{3}(x_{1},x_{2},x_{3})$, in case $\Delta=1^{3}$;
  2. $\bullet$
    $\Gamma_{3}(x_{1},x_{2},x_{1})$, in case $\Delta=(2,1)$,
  3. $\bullet$
    $\emptyset$, in case $\Delta=(3)$.
Proof.
1. It follows from Lemmas 4.1 and 4.2.
2. Lemma 6.2 implies that $\Gamma_{3}(x_{1},x_{2},x_{1})\neq 0$ is a basis for ${{\rm Id}({\mathsf{A}_{1},\mathsf{V}})}_{(2,1)}$ and $\emptyset$ is a basis for ${{\rm Id}({\mathsf{A}_{1},\mathsf{V}})}_{(2,1)}=\{0\}$.
Since $\Gamma_{3}(x_{1},x_{2},x_{3})$ and ${\rm St}_{3}(x_{1},x_{2},x_{3})$ are multilinear, Lemma 6.2 implies that every element $f\in{{\rm Id}({\mathsf{A}_{1},\mathsf{V}})}_{1^{3}}$ lies in the $\mathbb{F}$-span of
$\Gamma_{3}(x_{\sigma(1)},x_{\sigma(2)},x_{\sigma(3)})\;\text{ and }\;{\rm St}_{3}(x_{\sigma(1)},x_{\sigma(2)},x_{\sigma(3)})$
for all $\sigma\in S_{3}$. Since $\Gamma_{3}(x_{2},x_{3},x_{1})=-\Gamma_{3}(x_{1},x_{2},x_{3})+\Gamma_{3}(x_{1},x_{3},x_{2})$, we obtain that $f$ lies in the $\mathbb{F}$-span of $\Gamma_{3}(x_{1},x_{2},x_{3})$, $\Gamma_{3}(x_{1},x_{3},x_{2})$, and ${\rm St}_{3}(x_{1},x_{2},x_{3})$. The linear independence follows from straightforward calculations. ∎
7. Weak polynomial identities of degrees 4 and 5
Proposition 7.1.
Any weak polynomial identity for the pair $(\mathsf{A}_{1},\mathsf{V})$ of degree $4$ lies in the L-ideal $\mathcal{I}$ generated by $\Gamma_{3}$, ${\rm St}_{3}$, and $T_{4}$.
Proof.
Assume that $f\in\mathbb{F}\langle X\rangle$ is a weak polynomial identity of degree $4$ for the pair $(\mathsf{A}_{1},\mathsf{V})$. By Lemma 3.2, we can assume that $f$ is multihomogeneous of multidegree $\Delta$ with $|\Delta|=4$. By Theorem 5.10, $f$ is equivalent to a linear combination of completely reduced bracket-monomials of multidegree $\Delta$.
Assume $\Delta=1^{4}$. Using Example 5.4 we can see that
$\begin{array}[]{ccccccc}f(x_{1},\ldots,x_{4})&\equiv&\alpha_{1}\,x_{1}x_{2}[x_{3},x_{4}]&+&\alpha_{2}\,x_{1}x_{3}[x_{2},x_{4}]&+&\alpha_{3}\,x_{2}x_{3}[x_{1},x_{4}]\\ &+&\beta_{1}[x_{1},x_{2}][x_{3},x_{4}]&+&\beta_{2}[x_{1},x_{3}][x_{2},x_{4}],&\\ \end{array}$
where $\alpha_{i},\beta_{j}\in\mathbb{F}$. Since $0=f(x,x,y,x)=\alpha_{1}x^{2}$, $0=f(x,y,x,x)=-\alpha_{2}x^{2}$, and $0=f(y,x,x,x)=-\alpha_{3}x^{2}$, we obtain $\alpha_{1}=\alpha_{2}=\alpha_{3}=0$. Then equalities $0=f(x,y,x,y)=\beta_{1}$ and $0=f(x,x,y,y)=\beta_{2}$ imply that $f=0$.
Assume $\Delta=(2,1,1)$. Then
$f(x_{1},x_{2},x_{3})\equiv\alpha_{1}\,x_{1}^{2}[x_{2},x_{3}]+\alpha_{2}\,x_{1}x_{2}[x_{1},x_{3}]+\alpha_{3}\,[x_{1},x_{2}][x_{1},x_{3}],$
where $\alpha_{1},\alpha_{2},\alpha_{3}\in\mathbb{F}$. We have $\alpha_{1}=0$, since $0=f(x,y,x)=\alpha_{1}x^{2}$. Thus, the equality $0=f(x,x,y)=-\alpha_{2}x^{2}$ implies $\alpha_{2}=0$. Finally, since $0=f(y,x,x)=\alpha_{3}$, we obtain that $f=0$.
If $\Delta$ belongs to the list $\{(3,1),\;(2,2),\;(4)\}$, then Theorem 6.1 concludes the proof. ∎
Proposition 7.2.
Any weak polynomial identity for the pair $(\mathsf{A}_{1},\mathsf{V})$ of degree $5$ lies in the L-ideal $\mathcal{I}$ generated by $\Gamma_{3}$, ${\rm St}_{3}$, and $T_{4}$.
Proof.
Assume that $f\in\mathbb{F}\langle X\rangle$ is a weak polynomial identity of degree $5$ for the pair $(\mathsf{A}_{1},\mathsf{V})$. By Lemma 3.2, we can assume that $f$ is multihomogeneous of multidegree $\Delta$ with $|\Delta|=5$. By Theorem 5.10, $f$ is equivalent to a linear combination of completely reduced bracket-monomials of multidegree $\Delta$.
Assume $\Delta=1^{5}$. Using Example 5.4 we can see that $f(x_{1},\ldots,x_{5})$ is equivalent to
$\alpha_{1}\,x_{1}x_{2}x_{3}[x_{4},x_{5}]+\alpha_{2}\,x_{1}x_{2}x_{4}[x_{3},x_{5}]+\alpha_{3}\,x_{1}x_{3}x_{4}[x_{2},x_{5}]+\alpha_{4}\,x_{2}x_{3}x_{4}[x_{1},x_{5}]$
$+\beta_{1}\,x_{1}[x_{2},x_{3}][x_{4},x_{5}]+\beta_{2}\,x_{1}[x_{2},x_{4}][x_{3},x_{5}]+\beta_{3}\,x_{2}[x_{1},x_{3}][x_{4},x_{5}]+\beta_{4}\,x_{2}[x_{1},x_{4}][x_{3},x_{5}]$
$+\beta_{5}\,x_{3}[x_{1},x_{4}][x_{2},x_{5}]$
for some $\alpha_{i},\beta_{j}\in\mathbb{F}$. Considering $f(x,x,x,y,x)=f(x,x,y,x,x)=f(x,y,x,x,x)=f(y,x,x,x,x)=0$, we obtain $\alpha_{1}=\alpha_{2}=\alpha_{3}=\alpha_{4}=0$. Equalities $f(y,y,x,x,x)=f(y,x,y,x,x)=f(x,y,y,x,x)=f(y,x,x,y,x)=0$ imply that $\beta_{5}=\beta_{4}=\beta_{2}=\beta_{3}=0$. Finally, $0=f(x,y,x,y,x)=\beta_{1}x$ implies $\beta_{1}=0$, i.e., $f=0$.
Assume $\Delta=(3,1,1)$. Then
$f(x_{1},x_{2},x_{3})\equiv\alpha_{1}x_{1}^{3}[x_{2},x_{3}]+\alpha_{2}x_{1}^{2}x_{2}[x_{1},x_{3}]+\alpha_{3}x_{1}[x_{1},x_{2}][x_{1},x_{3}]$
for some $\alpha_{i}\in\mathbb{F}$. We have $\alpha_{1}=0$, since $0=f(x,y,x)=\alpha_{1}x^{3}$. Thus, the equality $0=f(x,x,y)=-\alpha_{2}x^{3}$ implies $\alpha_{2}=0$. Finally, since $0=f(y,x,x)=\alpha_{3}y$, we obtain that $f=0$.
Assume $\Delta=(2,2,1)$. Then $f(x_{1},x_{2},x_{3})$ is equivalent to
$\alpha_{1}x_{1}^{2}x_{2}[x_{2},x_{3}]+\alpha_{2}x_{1}x_{2}^{2}[x_{1},x_{3}]+\alpha_{3}x_{1}[x_{1},x_{2}][x_{2},x_{3}]+\alpha_{4}x_{2}[x_{1},x_{2}][x_{1},x_{3}]$
for some $\alpha_{i}\in\mathbb{F}$. We have $\alpha_{1}=\alpha_{3}=0$, since $0=f(x,y,x)=\alpha_{1}x^{2}y-\alpha_{3}x$. Thus, the equality $0=f(x,x,y)=-\alpha_{2}x^{3}$ implies $\alpha_{2}=0$. Finally, since $0=f(y,x,x)=\alpha_{4}x$, we obtain that $f=0$.
Assume $\Delta=(2,1,1,1)$. Then $f(x_{1},x_{2},x_{3},x_{4})$ is equivalent to
$\alpha_{1}x_{1}^{2}x_{2}[x_{3},x_{4}]+\alpha_{2}x_{1}^{2}x_{3}[x_{2},x_{4}]+\alpha_{3}x_{1}x_{2}x_{3}[x_{1},x_{4}]$
$\beta_{1}x_{1}[x_{1},x_{2}][x_{3},x_{4}]+\beta_{2}x_{1}[x_{1},x_{3}][x_{2},x_{4}]+\beta_{3}x_{2}[x_{1},x_{3}][x_{1},x_{4}]$
for some $\alpha_{i},\beta_{i}\in\mathbb{F}$. We have $\alpha_{1}=\alpha_{2}=0$, since $0=f(x,x,y,x)=\alpha_{1}x^{3}$, $0=f(x,y,x,x)=\alpha_{2}x^{3}$. Thus, the equality $0=f(x,x,x,y)=-\alpha_{3}x^{3}$ implies $\alpha_{3}=0$. Considering $0=f(y,x,x,x)=\beta_{3}x$, we obtain $\beta_{3}=0$. Finally, equalities $0=f(y,y,x,x)=\beta_{2}y$ and $0=f(y,x,y,x)=\beta_{1}y$ imply that $\beta_{1}=\beta_{2}=0$, i.e., $f=0$.
If $\Delta$ belongs to the list $\{(4,1),\;(3,2),\;(5)\}$, then Theorem 6.1 concludes the proof. ∎